OpenStudy (anonymous):

show that for all natural numbers, n 1. n^5 - n is divisible by 5. any help please?

OpenStudy (anonymous):

but that's not true... for example, if n=2, a natural number, then n^5=32. but 32 is NOT divisible by 5. can you check if the problem is written correctly?

OpenStudy (anonymous):

oops!!. mistake!! EDIT: n^5 - n is divisible by 5

OpenStudy (anonymous):

ok... maybe it'll help if we factor first: \(\large n^5-n=n(n^4-1)=n(n^2+1)(n^2-1)=n(n^2+1)(n+1)(n-1) \)

OpenStudy (anonymous):

i'm just thinking aloud but: notice that n, (n+1), (n-1) are three consecutive natural numbers...

OpenStudy (anonymous):

but if n = 4 then this will be come 4, 5, 3 which are not consecutive....

OpenStudy (anonymous):

well, not in that order that you wrote. but 3, 4, 5 are consecutive

OpenStudy (anonymous):

and when n=4, since one of the factors is a 5, that proves that when n=4, n^4-n is divisible by 5.

OpenStudy (anonymous):

so if n = 10 i would get 9, 10, 11 i haven't shown n^5 - n is divisible by 5 then :(

OpenStudy (anonymous):

yes... the proof has to be more formal than this but if n=10, 10 is divisible by 5 so 10^5-10 is

OpenStudy (anonymous):

*also

OpenStudy (anonymous):

we know this. but HOW do i show it? how do i make it "formal"?

OpenStudy (anonymous):

i'm sorry... i'm hitting dead ends with whatever i'm trying...

OpenStudy (anonymous):

is this for an analysis class?

OpenStudy (anonymous):

here's a proof by induction of the expression you want: http://www.physicsforums.com/showthread.php?t=245925

OpenStudy (anonymous):

thank you very much...